OpenStudy (anonymous):

Use the table of the function f(x) = x4 - 2x3 to answer this question. x f(x) -1 3 0 0 1 -1 2 0 3 27 What is the average rate of change from x = -1 to x = 1? -2 0 2 4

OpenStudy (dumbcow):

use slope formula\[slope = \frac{f(x_1) - f(x_0)}{x_1 - x_0}\] \[= \frac{-1 - 3}{1 - (-1)}\]

OpenStudy (anonymous):

-2